Services Status: 3DOWN

IS&T Home > 3-Down

What you’ll find on this page:
General Services
Network
E-mail
Telephone Service
Academic Services
Library Services
Administrative Services

Greetings. Today is Thursday, February 23, 2017. This "3DOWN" Services Status page lists the status of the major services available to the MIT community, including scheduled and unscheduled outages.

We have had 8538 hits today with 8498 unique visitors.

General Services
Last Update: Thursday, February 23, 2017 07:05 AM

Feburary 23, 2017 7:01am.  The network issue impacting servers in the OC11 data center from 5:56am to 7:00am has been resolved.  

 

The Tech Shuttle tracking issue within the MIT Mobile app and website has been resolved.

February 18, 2017 03:00PM: Planned SAP Production Outage on Saturday, February 18. SAP and IS&T have completed this weekend's disaster recovery testing. The SAP production environment was unavailable on Saturday, February 18th, between 04:00AM and around 12:15PM. For the duration of the exercise, SAP, ECAT, and all Atlas SAP activity, including time sheet entry/approval, training, requisitioning, journal vouchers, credit card verification, events planning and service requests, were unavailable to the MIT community. SAP production services are now available in accordance with their normal schedules.

All services are operating normally.

 

February 28th, 2017 2:00pm: Planned maintenance of TSM server oc11-bk-ent-1.  IS&T needs to perform some maintenance of the server oc11-bk-ent-1 at 2pm on the 28th.  The server will be unavailable for a short period of time while it is rebooted to remove some old devices left over from a migration.  We expect the activity to take between 15 and 30 minutes.

Back To Top
Network
Last Update: Wednesday, February 22, 2017 07:03 AM

All services are operating normally.

 

2/22/17 6:00AM - 7:30AM EST [COMPLETED]:  IS&T will be moving the APs in NW10 to a dedicated AP Group. A brief 5-10 min outage is expected for each AP in NW10 during the maintenance window. 

Back To Top
Email
Last Update: Monday, January 23, 2017 10:47 AM

1/23/2017 10:30AM: IS&T has resolved the issue with the Exchange service. We apologize for the incoveinence.

Back To Top
Telephone Service
Last Update: Monday, February 20, 2017 07:38 AM

All services are operating normally.

Back To Top
Academic Services
Last Update: Sunday, February 19, 2017 07:21 AM

All services are operating normally.

Back To Top
Athena Message of the Day
Last Update: Friday, September 30, 2016 07:56 PM
On October 8, 2016 the 38-370 Athena cluster will be closed permanently. The Athena cluster combo will be changed on October 19, 2016. Please run: tellme combo at an Athena prompt to see the new combo.
Back To Top
Library Services
Last Update: Monday, September 26, 2016 12:57 PM

We are aware of a problem with the display of MIT-specific links in Google Scholar

ScienceDirect (Elsevier): Access appears to be restored. If you continue to experience problems, please contact us at Ask Us .
Back To Top
Administrative Services
Last Update: Saturday, February 18, 2017 03:01 PM

IS&T is aware of problems accessing citrixapps.mit.edu. Please use the URL https://citrix.mit.edu  to access the new site and remove any existing bookmarks you may have to the old URL.

February 18, 2017 03:00PM: Planned SAP Production Outage on Saturday, February 18. SAP and IS&T have completed this weekend's disaster recovery testing. The SAP production environment was unavailable on Saturday, February 18th, between 04:00AM and around 12:15PM. For the duration of the exercise, SAP, ECAT, and all Atlas SAP activity, including time sheet entry/approval, training, requisitioning, journal vouchers, credit card verification, events planning and service requests, were unavailable to the MIT community. SAP production services are now available in accordance with their normal schedules.

All services are operating normally.

Previous Versions

Follow the link below for the previous version of the Services Status page.
Thursday, February 23, 2017 7:05 AM